# Scaling limit of heavy-tailed nearly unstable cumulative INAR($\infty$) processes and rough fractional diffusions

### v6

@@ 2 Assumptions and Main Results
$\rho^{T}(x):=T\frac{A_{\lfloor xT\rfloor}^{T}}{\left\lVert A^{T}\right\rVert_{1}}.$ (10)
Each $\rho^{T}$ is a step function and a valid probability density function, since it is non-negative and integrates to one. As $T\to\infty$, the sequence of probability measures corresponding to these densities converges weakly to a probability measure on $0,\infty)$ with the density $\lambda x^{\alpha-1}E_{\alpha,\alpha}(-\lambda x^{\alpha})$.
Furthermore, let $F^{T}(t)$ be the associated cumulative distribution function, defined by $F^{T}(t)=\int_{0}^{t}\rho^{T}(x)dx$. Then, over the compact interval $[0,1]$, $F^{T}(t)$ converges uniformly to the limiting distribution function $F^{\alpha,\lambda}(t)$ defined in (9).
@@
We provide the proof in Section 4.1. ∎
To prove this lemma, we need a discrete-type extended Karamata-Tauberian Theorem, which is similar to the Theorem 1.7.6 in Bingham et al. (1989). We first give the discrete versions of Proposition 1.5.8, Theorem 1.7.6 in Bingham et al. (1989).
Lemma 2.
Let $l(t)$ be a slowly varying function that is locally bounded on $X_{0},\infty)$ for some $X_{0}>0$, and let $\alpha>-1$. For any fixed integer $X\geq X_{0}$, the following holds for integers $x\to\infty$:
$\sum_{t=X}^{x}t^{\alpha}l(t)\sim\frac{1}{\alpha+1}x^{\alpha+1}l(x).$
Proof.
We provide the proof in Section 4.2. ∎
Lemma 3 (Discrete-type extended Karamata-Tauberian Theorem).
@@
We provide the proof in Section 4.3. ∎
The tightness of $(Z^{T},Y^{T})$ also requires proof. Therefore, we start with the following lemma.
Lemma 4.
@@
We provide the proof in Section 4.4. ∎
Since $(Y^{T})$ and $(\Lambda^{T})$ are pure jump processes, the classical Kolmogorov-Chentsov tightness criterion is not applicable. To establish the tightness of $(Z^{T},Y^{T})$, we instead apply a recent criterion from Horst et al. (2023), which we restate as Lemma 5 below for the reader’s convenience. This result gives a new criterion to verify the $C$-tightness for a sequence of càdlàg processes and will be used in proving Lemma 4. Additionally, proving the convergence of the sharp bracket of $Z^{T}$ in the discrete-time case presents more challenges than in the continuous-time case discussed by Jaisson and Rosenbaum (2016). For instance, in a Hawkes process $(N_{t})_{t\geq 0}$ with intensity function is
@@
the angle bracket of $(N_{t})_{t\geq 0}$ is given by $\int_{0}^{\cdot}\lambda_{s}\,ds$, and the sharp bracket is simply $N_{\cdot}$, which is straightforward to derive. However, achieving a similar result in the discrete-time setting is not as immediate and requires more nuanced handling. Indeed, the value $\lambda_{n}$ in (2) is analogous the intensity process $\lambda_{t}$ in continuous-time Hawkes processes. In the proof section, we work directly with $(Y^{T})$, and in Section 2.1 provides an intuitive interpretation of the rescaled $\lambda^{T}$. To clarify the proof clear, we introduce following lemma here.
Lemma 5 (Criterion for $C$-tightness, Lemma 3.5 in Horst et al. (2023)).
@@
Then the sequence $(X^{(n)})_{n\geq 1}$ is $C$-tight on $[0,T]$ (i.e., tight in the space of continuous functions $C([0,T])$ equipped with the uniform topology).
Additionally, we need the following lemma to prove tightness and convergence in the Skorohod topology.
Lemma 6.
The sequence of martingales $Y^{T}-\Lambda^{T}$ tends to zero in probability, uniformly on $[0,1]$.
@@
We provide the proof in Section 4.5. ∎
Now we can give the main result.
Proposition 1 (Tightness of the Sequence).
Under Assumption 1 and 2, the sequence $(Z^{T},Y^{T})$ is tight. Furthermore, if $(Z,Y)$ is a limit point of $(Z^{T},Y^{T})$ as $T\rightarrow\infty$, then $Z$ is a continuous martingale and $[Z,Z]=Y$.
@@
We provide the proof in Section 5.1. ∎
Next, we characterize the limit process $Y$ (see Theorem 1). To this end, we first introduce a lemma that solves discrete renewal equations, which will be essential for proving Theorem 1.
Lemma 7 (Solution of Discrete Renewal Equations).
@@
We provide the proof in Section 4.6. ∎
Theorem 1 (Characterization of the Limit Process).
For any limit point $(Z,Y)$ defined in Proposition 1, there exists a limiting probability space $(\Omega,\mathcal{F},\mathbb{P})$ on which the convergence holds almost surely. On this probability space, there exists a Brownian motion $B$ (up to extension) such that for all $t\in[0,1]$, $Z_{t}=B_{Y_{t}}$. Furthermore, for any $\epsilon>0$, $Y$ is continuous with Hölder regularity $(1\wedge 2\alpha)-\epsilon$ on $[0,1]$ and satisfies
$Y_{t}=\int_{0}^{t}sf^{\alpha,\lambda}(t-s)\,ds+\frac{1}{\sqrt{\nu^{*}\lambda}}\int_{0}^{t}f^{\alpha,\lambda}(t-s)B_{Y_{s}}\,ds.$ (12)
@@
The proof of the Hölder property for $Y$ in Theorem 1 follows exactly the same argument as the Proof of Theorem 3.1 in Jaisson and Rosenbaum (2016). The derivation of the equation (12) is provided in Section 5.2. ∎
After (12) in Theorem 1 is established, we can characterize its derivative process as follows. This result is from Theorem 3.2 in Jaisson and Rosenbaum (2016).
Corollary 1 (Characterization of the Derivative Process).
Let $(Y_{t})$ be a process satisfying (12) for $t\in[0,1]$ and assume $\alpha>1/2$. Then $Y$ is differentiable on $[0,1]$ and its derivative $\dot{Y}$ satisfies
@@
where $B$ is the same Brownian motion as in (12). Furthermore, for any $\epsilon>0$, $\dot{Y}$ has Hölder regularity $\alpha-1/2-\epsilon$.
@@
The proof is exactly the same as the Proof of Theorem 3.2 in Jaisson and Rosenbaum (2016). ∎
@@ 2.1 Heuristic Motivation for the Scaling Limit
In the “light-tailed” case, when the condition $\sum_{n=1}^{\infty}n\eta_{n}<\infty$ is met as in Barczy et al. (2011), the scaling limit of the INAR process is known to be a squared Bessel process. Our work, however, focuses on the heavy-tailed case governed by Assumption 1, which induces a different scaling behavior.
@@
$C_{t}^{T}:=\frac{1-a_{T}}{\mu^{T}}\lambda_{\lfloor tT\rfloor}^{T}.$ (13)
@@
$\lambda_{n}^{T}=\mu^{T}+\mu^{T}\sum_{s=1}^{n-1}A_{n-s}^{T}+\sum_{s=1}^{n-1}A_{n-s}^{T}\left(M_{s}^{T}-M_{s-1}^{T}\right).$ (14)
@@
We provide the proof in Section 4.7. ∎
We substitute the expression for $\lambda_{\lfloor tT\rfloor}^{T}$ from Lemma 8 to expand the definition of $C_{t}^{T}$:
$C_{t}^{T}=$ $\frac{(1-a_{T})}{\mu^{T}}\lambda_{\lfloor tT\rfloor}^{T}=\frac{(1-a_{T})}{\mu^{T}}\left(\mu^{T}+\mu^{T}\sum_{s=1}^{\lfloor tT\rfloor-1}A_{s}^{T}+\sum_{s=1}^{\lfloor tT\rfloor-1}A_{\lfloor tT\rfloor-s}^{T}(M_{s}^{T}-M_{s-1}^{T})\right)$ (15)
@@
The asymptotic form of the renewal sequence $A_{k}^{T}$ was established in the proof of Lemma 1. For large $T$ and $k$, we have the approximation:
$A_{k}^{T}\approx\frac{\left\lVert A^{T}\right\rVert_{1}v_{T}}{T}\left(\frac{k}{T}\right)^{\alpha-1}E_{\alpha,\alpha}\left(-v_{T}\left(\frac{k}{T}\right)^{\alpha}\right).$
Using this approximation, the sum can be viewed as a Riemann sum, which for large $T$ converges to the corresponding integral:
$\sum_{s=1}^{\lfloor tT\rfloor-1}A_{s}^{T}$ $=\sum_{s=1}^{\lfloor tT\rfloor-1}\frac{1}{T}\cdot\left(TA_{s}^{T}\right)$
$\sim\int_{0}^{t}TA_{\lfloor Tx\rfloor}^{T}dx$
$\sim\int_{0}^{t}\left\lVert A^{T}\right\rVert_{1}v_{T}x^{\alpha-1}E_{\alpha,\alpha}(-v_{T}x^{\alpha})dx,$
where $v_{T}$ is defined in (23). Define
@@
$(M_{s}^{T}-M_{s-1}^{T})_{s\geq 1}$ is a martingale difference sequence, and it is obvious that
$\sum_{s=1}^{\lfloor tT\rfloor-1}\mathbb{E}\left[\frac{1}{T}\frac{(M_{s}^{T}-M_{s-1}^{T})^{2}}{\lambda_{s}^{T}}\right]=$ $\frac{\lfloor tT\rfloor-1}{T}\rightarrow t,\ \text{as}\ T\rightarrow\infty.$
We also need to check the conditional Lindeberg condition for the martingale central limit theorem. Let the martingale differences be $\Delta B_{s}^{T}=\frac{1}{\sqrt{T}}\frac{X_{s}^{T}-\lambda_{s}^{T}}{\sqrt{\lambda_{s}^{T}}}$. The condition requires that for any $\epsilon>0$, the sum of conditional expectations,
$L_{T}:=\sum_{s=1}^{\lfloor tT\rfloor}\mathbb{E}\left[(\Delta B_{s}^{T})^{2}\cdot 1_{\{|\Delta B_{s}^{T}|>\epsilon\}}\bigg|\mathcal{F}_{s-1}\right],$
converges to 0 in probability as $T\to\infty$.
To prove this, we first find a bound for each term in the sum using a standard moment inequality. Given the history $\mathcal{F}_{s-1}$, $\lambda_{s}^{T}$ is a known quantity. For the random variable $\Delta B_{s}^{T}$ and constant $\epsilon>0$, we have:
$\mathbb{E}\left[(\Delta B_{s}^{T})^{2}\cdot 1_{\{|\Delta B_{s}^{T}|>\epsilon\}}\bigg|\mathcal{F}_{s-1}\right]$ $\leq\frac{\mathbb{E}\left[(\Delta B_{s}^{T})^{4}\bigg|\mathcal{F}_{s-1}\right]}{\epsilon^{2}}$
$=\frac{1}{\epsilon^{2}T^{2}}\mathbb{E}\left[\left.\frac{(X_{s}^{T}-\lambda_{s}^{T})^{4}}{(\lambda_{s}^{T})^{2}}\right|\mathcal{F}_{s-1}\right]$
[…]
$=\frac{1}{\epsilon^{2}T^{2}}\left(\frac{1}{\lambda_{s}^{T}}+3\right).$
Summing this bound from $s=1$ to $\lfloor tT\rfloor$ provides a random upper bound for $L_{T}$:
$0\leq L_{T}\leq\frac{1}{\epsilon^{2}T^{2}}\sum_{s=1}^{\lfloor tT\rfloor}\left(\frac{1}{\lambda_{s}^{T}}+3\right)=:R_{T}.$
To show that $L_{T}\to 0$ in probability, it is sufficient to show that $R_{T}\to 0$ in probability. We prove this by showing that its expectation converges to zero, i.e., $\mathbb{E}[R_{T}]\to 0$.
$\mathbb{E}[R_{T}]$ $=\mathbb{E}\left[\frac{1}{\epsilon^{2}T^{2}}\sum_{s=1}^{\lfloor tT\rfloor}\left(\frac{1}{\lambda_{s}^{T}}+3\right)\right]=\frac{1}{\epsilon^{2}T^{2}}\sum_{s=1}^{\lfloor tT\rfloor}\left(3+\mathbb{E}\left[\frac{1}{\lambda_{s}^{T}}\right]\right).$
Since the process intensity is always bounded below by the baseline rate, $\lambda_{s}^{T}\geq\mu^{T}>0$, we have a bound for the expectation:
$\mathbb{E}\left[\frac{1}{\lambda_{s}^{T}}\right]\leq\frac{1}{\mu^{T}}.$
Substituting this into the expression for $\mathbb{E}[R_{T}]$:
$\mathbb{E}[R_{T}]$ $\leq\frac{1}{\epsilon^{2}T^{2}}\sum_{s=1}^{\lfloor tT\rfloor}\left(3+\frac{1}{\mu^{T}}\right)=\frac{\lfloor tT\rfloor}{\epsilon^{2}T^{2}}\left(3+\frac{1}{\mu^{T}}\right).$
From Assumption 2, we know $\mu^{T}\sim CT^{\alpha-1}$ for some constant $C>0$. Since $\alpha>1/2$, we have $\mathbb{E}[R_{T}]\to 0$ as $T\to\infty$. As $R_{T}$ is a non-negative random variable, convergence of its expectation to zero implies that $R_{T}\to 0$ in probability. As a result, $L_{T}$ must also converge to 0 in probability, which rigorously verifies the conditional Lindeberg condition.
By the martingale central limit theorem (see Section 3 in Hall and Heyde (1980)), $B_{t}^{T}$ converges to a standard Brownian motion. For sufficiently large $T$, from Lemma 1,
$\sum_{s=1}^{\lfloor tT\rfloor-1}A_{\lfloor tT\rfloor-s}^{T}(M_{s}^{T}-M_{s-1}^{T})$ (16)
$=$ $\sum_{s=1}^{\lfloor tT\rfloor-1}A_{\lfloor tT-s\rfloor}^{T}(M_{s}^{T}-M_{s-1}^{T})$
$=$ $\sum_{s=1}^{\lfloor tT\rfloor-1}A_{\lfloor T\left(t-\frac{s}{T}\right)\rfloor}^{T}(M_{s}^{T}-M_{s-1}^{T})$
$\approx$ $\sum_{s=1}^{\lfloor tT\rfloor-1}\frac{\left\lVert A^{T}\right\rVert_{1}}{T}v_{T}\cdot\left(t-\frac{s}{T}\right)^{\alpha-1}E_{\alpha,\alpha}\left(-v_{T}\left(t-\frac{s}{T}\right)^{\alpha}\right)\sqrt{T\lambda_{s}^{T}}\left(B_{(s+1)/T}^{T}-B_{s/T}^{T}\right).$
So here we face the same problem as in Jaisson and Rosenbaum (2016): to ensure that the limits of $v_{T}$ and another related term $\gamma_{T}$ are positive constants. This motivates our Assumption 2, which is consistent with the framework in Jaisson and Rosenbaum (2016).
Under Assumption 2, let
$\kappa_{s}^{T}=\sum_{u=1}^{s-1}A_{s-u}^{T}(M_{u}^{T}-M_{u-1}^{T}),$

### v7

@@ 2.2 A key discrete Laplace limit and tightness
$\rho^{T}(x):=T\frac{A_{\lfloor xT\rfloor}^{T}}{\left\lVert A^{T}\right\rVert_{1}}.$ (2.8)
Let $P^{T}$ be the probability measure on $0,\infty)$ with density $\rho^{T}$, and let $F^{T}(t):=\int_{0}^{t}\rho^{T}(x)\,dx$ be its distribution function. Then, as $T\to\infty$, $P^{T}$ converges weakly to the probability measure $P$ on $0,\infty)$ with density $f^{\alpha,\lambda}$ defined in (2.5). Moreover,
$\sup_{t\in[0,1]}\left|F^{T}(t)-F^{\alpha,\lambda}(t)\right|\to 0,$}
where $F^{\alpha,\lambda}$ is defined in (2.6).
@@
We provide the proof in Section A.2. ∎
To prove Lemma 2.4, we will use the following discrete-type extended Karamata–Tauberian theorem.
Lemma 2.5 (Discrete-type extended Karamata-Tauberian Theorem).
@@
We provide the proof in Section A.3. ∎
We now record the tightness inputs.
Lemma 2.6.
@@
We provide the proof in Section A.4. ∎
In the tightness proof we will repeatedly bound expressions involving the resolvent coefficients, in particular sums of the form $\sum_{k\leq s}(A_{k}^{T})^{2}$. While Lemma 2.4 provides a Laplace/weak-convergence characterization of the rescaled resolvent measure (useful later for identifying the limit), it does not imply pointwise bounds for $A_{k}^{T}$. We therefore establish below a deterministic polynomial upper bound, uniform in $T$, which will be the only pointwise input needed in the tightness argument.
Lemma 2.7 (Uniform polynomial bound for the resolvent).
Let $(\eta_{n})_{n\geq 1}$ satisfy Assumption 2.1(i) with $\alpha\in(1/2,1)$ and $\sum_{n\geq 1}\eta_{n}=1$. Define the (critical) renewal/resolvent sequence
$A_{k}:=\sum_{m=1}^{\infty}(\eta^{*m})_{k},\qquad k\geq 1.$
Then there exists a constant $C>0$ such that for all $k\geq 1$,
$A_{k}\leq C\,k^{\alpha-1}.$
Consequently, for every $T$ and every $k\geq 1$,
$A_{k}^{T}=\sum_{m=1}^{\infty}(\eta^{T*m})_{k}=\sum_{m=1}^{\infty}a_{T}^{m}(\eta^{*m})_{k}\leq A_{k}\leq C\,k^{\alpha-1}.$
Proof.
We provide the proof in Section A.5. ∎
Since $(Y^{T})$ and $(\Lambda^{T})$ are pure jump processes, the classical Kolmogorov-Chentsov tightness criterion is not applicable. To establish the tightness of $(Z^{T},Y^{T})$, we instead apply a recent criterion from Horst et al. (2023), which we restate as Lemma 2.8 below for the reader’s convenience. This result gives a new criterion to verify the CC–tightness for a sequence of càdlàg processes and will be used in proving Lemma 2.6. Additionally, proving the convergence of the sharp bracket of $Z^{T}$ in the discrete-time case presents more challenges than in the continuous-time case discussed by Jaisson and Rosenbaum (2016). For instance, in a Hawkes process $(N_{t})_{t\geq 0}$ with intensity function
@@
the angle bracket of $(N_{t})_{t\geq 0}$ is given by $\int_{0}^{\cdot}\lambda_{s}\,ds$, and the sharp bracket is simply $N_{\cdot}$, which is straightforward to derive. However, achieving a similar result in the discrete-time setting is not as immediate and requires more nuanced handling. Indeed, the value $\lambda_{n}$ in (1.2) is analogous to the intensity process $\lambda_{t}$ in continuous-time Hawkes processes. In the proof section, we work directly with $(Y^{T})$, and Section 2.3 provides an intuitive interpretation of the rescaled $\lambda^{T}$. To make the proof clear, we recall the following result.
Lemma 2.8 (Criterion for $C$-tightness, Lemma 3.5 in Horst et al. (2023)).
@@
Then the sequence $(X^{(n)})_{n\geq 1}$ is $C$-tight on $[0,T]$.
Lemma 2.9.
The martingale part of the rescaled counting process vanishes: as $T\to\infty$,
$\sup_{t\in[0,1]}\left|Y_{t}^{T}-\Lambda_{t}^{T}\right|\xrightarrow{\mathbb{P}}0.$}
Equivalently, $Y^{T}-\Lambda^{T}$ converges to 0 in probability uniformly on $[0,1]$.
@@
We provide the proof in Section A.6. ∎
We now state the tightness and quadratic variation result for the sequence $(Z^{T},Y^{T})_{T\geq 1}$.
Proposition 2.10 (Tightness of the Sequence).
Under Assumption 2.1 and 2.3, the sequence $(Z^{T},Y^{T})$ is tight. Furthermore, if $(Z,Y)$ is a limit point of $(Z^{T},Y^{T})$ as $T\rightarrow\infty$, then $Z$ is a continuous martingale and $[Z,Z]=Y$.
@@
We provide the proof in Section B.1. ∎
Renewal equation and characterization of the limit.
To characterize the limit $Y$, we will repeatedly use the explicit solution of discrete renewal equations.
Lemma 2.11 (Solution of Discrete Renewal Equations).
@@
We provide the proof in Section A.7. ∎
Remark 2.12 (Mean intensity and the renewal sequence).
Recall that $\lambda_{s}^{T}=\mu^{T}+\sum_{k=1}^{s-1}\eta_{k}^{T}X_{s-k}^{T}$. Taking expectations and using $\mathbb{E}[X_{s-k}^{T}]=\mathbb{E}[\lambda_{s-k}^{T}]$ yields the renewal-type equation
$q_{s}^{T}:=\mathbb{E}[\lambda_{s}^{T}]=\mu^{T}+\sum_{k=1}^{s-1}\eta_{k}^{T}q_{s-k}^{T},\qquad s\geq 1.$
By Lemma 2.11, this admits the explicit solution
$q_{s}^{T}=\mu^{T}+\mu^{T}\sum_{u=1}^{s-1}A_{u}^{T},\qquad s\geq 1,$
where $(A_{u}^{T})_{u\geq 1}$ is defined in (2.1).
Theorem 2.13 (Characterization of the Limit Process).
For any limit point $(Z,Y)$ defined in Proposition 2.10, there exists a Brownian motion $B$ (possibly on an extension of the underlying probability space) such that for all $t\in[0,1]$, $Z_{t}=B_{Y_{t}}$. Furthermore, for any $\epsilon>0$, $Y$ is continuous with Hölder regularity $(1\wedge 2\alpha)-\epsilon$ on $[0,1]$ and satisfies
$Y_{t}=\int_{0}^{t}sf^{\alpha,\lambda}(t-s)\,ds+\frac{1}{\sqrt{\nu^{*}\lambda}}\int_{0}^{t}f^{\alpha,\lambda}(t-s)B_{Y_{s}}\,ds.$ (2.10)
@@
The proof of the Hölder property for $Y$ in Theorem 2.13 follows exactly the same argument as the Proof of Theorem 3.1 in Jaisson and Rosenbaum (2016). The derivation of the equation (2.10) is provided in Section B.2. ∎
After (2.10) is established, we can characterize the derivative process (instantaneous variance in the financial interpretation).
Corollary 2.14 (Characterization of the Derivative Process).
Let $(Y_{t})$ be a process satisfying (2.10) for $t\in[0,1]$ and assume $\alpha>1/2$. Then $Y$ is differentiable on $[0,1]$ and its derivative $\dot{Y}$ satisfies
@@
where $B$ is the same Brownian motion as in (2.10). Furthermore, for any $\epsilon>0$, $\dot{Y}$ has Hölder regularity $\alpha-1/2-\epsilon$.
@@
The proof is exactly the same as the Proof of Theorem 3.2 in Jaisson and Rosenbaum (2016). More details on the proof are given in Section B.3. ∎
@@ 2.3 Heuristic Motivation for the Scaling Limit
In the “light-tailed” case, when the condition $\sum_{n=1}^{\infty}n\eta_{n}<\infty$ is met as in Barczy et al. (2011), the scaling limit of the INAR process is known to be a squared Bessel process. Our work, however, focuses on the heavy-tailed case governed by Assumption 2.1, which induces a different scaling behavior.
@@
$C_{t}^{T}:=\frac{1-a_{T}}{\mu^{T}}\lambda_{\lfloor tT\rfloor}^{T}.$ (2.11)
@@
$\lambda_{n}^{T}=\mu^{T}+\mu^{T}\sum_{s=1}^{n-1}A_{n-s}^{T}+\sum_{s=1}^{n-1}A_{n-s}^{T}\left(M_{s}^{T}-M_{s-1}^{T}\right).$ (2.12)
@@
We provide the proof in Section A.8. ∎
We substitute (2.12) into (2.11) and obtain
$C_{t}^{T}=$ $\frac{(1-a_{T})}{\mu^{T}}\lambda_{\lfloor tT\rfloor}^{T}$ (2.13)
@@
The argument below is intended as intuition only. In particular, Lemma 2.4 provides convergence of Laplace transforms (i.e. weak convergence at the level of measures), which does not imply convergence of densities or pointwise asymptotics for $A_{k}^{T}$. Nevertheless, guided by the Laplace-transform asymptotics in Lemma 2.4, one may formally expect that for large $T$ and $k$ with $k/T$ of order one,
$A_{k}^{T}\approx\frac{\|A^{T}\|_{1}v_{T}}{T}\left(\frac{k}{T}\right)^{\alpha-1}E_{\alpha,\alpha}\!\left(-v_{T}\left(\frac{k}{T}\right)^{\alpha}\right).$
With this heuristic approximation in mind, the sum may be viewed as a Riemann sum:
$\sum_{s=1}^{\lfloor tT\rfloor-1}A_{s}^{T}$ $=\sum_{s=1}^{\lfloor tT\rfloor-1}\frac{1}{T}\cdot(TA_{s}^{T})\approx\int_{0}^{t}TA_{\lfloor Tx\rfloor}^{T}\,dx\approx\int_{0}^{t}\|A^{T}\|_{1}v_{T}x^{\alpha-1}E_{\alpha,\alpha}(-v_{T}x^{\alpha})\,dx.$
A fully rigorous treatment used later in the paper relies only on uniform bounds and weak convergence of the associated discrete measures (rather than density convergence); see the proof of the moment bounds in Section A.4.
Define
@@
By the martingale central limit theorem (see Section 3 in Hall and Heyde (1980)), $B_{t}^{T}$ converges to a standard Brownian motion. The last term in (2.13) can then be interpreted as a discrete convolution against a singular kernel, which motivates the limiting Volterra structure in Theorem 2.13.
Microstructural foundation for the initial value.
In practical applications, particularly in financial models such as the rough Heston model, it is essential to be able to prescribe a strictly positive initial instantaneous variance, i.e. $\dot{Y}_{0}=v_{0}$. In the classical Heston stochastic volatility model (see Heston (1993)), the asset price is driven by an instantaneous variance process $(V_{t})_{t\geq 0}$, and the integrated variance $Y_{t}:=\int_{0}^{t}V_{s}\,ds$ plays a central role in option pricing and likelihood-based inference. The rough Heston model (see El Euch and Rosenbaum (2019)) keeps the same interpretation—$V$ as instantaneous variance and $Y$ as integrated variance—but replaces the Markovian mean-reversion in $V$ by a fractional/Volterra-type kernel, yielding a non-Markovian variance process. In our scaling limit, the process $Y$ corresponds to the integrated variance, so that $\dot{Y}$ represents the instantaneous variance. A subtle point is that if one starts the limiting Volterra dynamics from a time-independent initial condition, the resulting trajectories satisfy $\dot{Y}_{0}=0$, which would prevent us from matching a given market/empirical level $v_{0}>0$. For this reason, we adopt a time-dependent initialization (equivalently, we add a suitable deterministic term at the origin) so that the limiting model allows an arbitrary prescribed value $\dot{Y}_{0}=v_{0}$.
Following the methodology of El Euch and Rosenbaum (2019) and its discrete-time adaptation in Wang et al. (2025), we introduce a time-dependent baseline intensity $\hat{\mu}_{T}(n)$.
We redefine the intensity as
$\lambda_{n}^{T}=\hat{\mu}_{T}(n)+\sum_{s=1}^{n-1}\eta_{n-s}^{T}X_{s}^{T},\quad n\in\mathbb{N},$ (2.14)
where the time-dependent baseline intensity is given by:
Definition 2.16 (Time-Dependent Baseline for Initial Value).
Fix $\xi\geq 0$. Define the time-dependent baseline intensity by
$\hat{\mu}_{T}(n):=\mu^{T}+\xi\mu^{T}\left(\frac{1}{1-a_{T}}\left(1-\sum_{s=1}^{n-1}\eta_{s}^{T}\right)-\sum_{s=1}^{n-1}\eta_{s}^{T}\right),\qquad n\in\mathbb{N},$
where $(\eta_{n}^{T})_{n\geq 1}$ is a (possibly $T$-dependent) kernel sequence satisfying $\sum_{n=1}^{\infty}\eta_{n}^{T}=a_{T}\in(0,1)$.
This baseline is strictly positive. Indeed, letting $S_{n-1}^{T}:=\sum_{s=1}^{n-1}\eta_{s}^{T}$, we have $0\leq S_{n-1}^{T}\leq a_{T}$, hence
$\frac{1}{1-a_{T}}(1-S_{n-1}^{T})-S_{n-1}^{T}\geq\frac{1}{1-a_{T}}(1-a_{T})-a_{T}=1-a_{T}>0,$
and therefore
$\hat{\mu}_{T}(n)\geq\mu^{T}+\xi\mu^{T}(1-a_{T})>0,\qquad\forall\,n\geq 1.$
Heuristic. At the very beginning ($n=1$), the sums are empty and
$\hat{\mu}_{T}(1)=\mu^{T}\left(1+\frac{\xi}{1-a_{T}}\right),$
which is large when $(1-a_{T})^{-1}$ is of order $\mathcal{O}(T^{\alpha})$, creating an initial burst of events. As $n$ increases, $S_{n-1}^{T}:=\sum_{s=1}^{n-1}\eta_{s}^{T}$ increases towards $a_{T}$, and the bracketed term
$\frac{1}{1-a_{T}}(1-S_{n-1}^{T})-S_{n-1}^{T}$
decreases from $\frac{1}{1-a_{T}}$ down to $1-a_{T}$. Consequently, $\hat{\mu}_{T}(n)$ relaxes towards the positive asymptotic level
$\lim_{n\to\infty}\hat{\mu}_{T}(n)=\mu^{T}+\xi\mu^{T}(1-a_{T})=\mu^{T}\left(1+\xi(1-a_{T})\right),$
rather than $\mu^{T}$. In the scaling regime i.e. $(1-a_{T})\to 0$, so this elevation is typically of lower order and $\hat{\mu}_{T}(n)$ may be viewed as approaching $\mu^{T}$ up to a vanishing correction. This engineered transient is what produces a non-zero initial value for the limiting rate process $\dot{Y}_{t}$.
With this modification, Theorem 2.13 extends to Corollary 2.18 below, showing that the limit $\dot{Y}_{t}$ starts from a value determined by $\xi$. We first introduce a lemma that will be used in the proof of Corollary 2.18.
Lemma 2.17 (Renewal Expansion with Time-Dependent Baseline).
Under the time-dependent baseline intensity $\hat{\mu}_{T}(n)$ from Definition 2.16, the cumulative intensity and the centered cumulative process have the following representations for $n\geq 1$:
$\mathbb{E}[N_{n}^{T}]=\sum_{i=1}^{n}\left(\hat{\mu}_{T}(i)+\sum_{s=1}^{i-1}A_{i-s}^{T}\hat{\mu}_{T}(s)\right),$ (2.15)